# FakeScope: Large Multimodal Expert Model for Transparent AI-Generated Image Forensics

### v1

@@ III-C1 Task Definition
FakeChain encapsulates forensic knowledge to enable models a fundamental understanding between visual trace evidence and authenticity. We structure it as a materialized reasoning process, comprising an exhaustive description of the appearance and location of trace evidence, enclosed with an authenticity conclusion. Compared with free-form descriptions, this structured reasoning format has two advantages: (1) it allows for a precise and detailed reflection of the aspects humans focus on when assessing authenticity; and (2) it inherently follows the logical pattern of human thought. For instance, when examining a suspicious image, humans first scrutinize telltale visual clues and then summarize an inclined conclusion like “Thus, I tend to think this is a fake image generated by AI”. This structured reasoning approach helps LMMs more effectively mimic human decision-making in AI-generated image detection, thereby enhancing human-centered transparency.
Figure 4: The composition of the FakeInstruct which is derived from the FakeChain dataset, containing 47K visual instructions on absolute authenticity judgments, 95K instructions on holistic reasoning, 715K on fine-grained visual trace evidence, and 1190K on extended knowledge and discussions. The million-scale diversified visual instructions of FakeInstruct enable LMMs with both broad and fine-grained understanding of image authenticity, ensuring LMMs are equipped to handle diverse forensic tasks with both interpretability and accuracy.
@@ III-C2 ACoTI Strategy
To develop a cost-efficient approach for obtaining large-scale reasoning data, we present a semi-automatic scheme to efficiently extract forensic knowledge from AI experts with few human annotations as supervision, i.e., the Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy. As illustrated in Fig. 3, ACoTI operates in a human-machine collaborative framework, where humans partially participate in the first and second phases by providing partial information of trace evidence. Inspired by human reasoning processes, ACoTI follows the Steer-Demonstrate-Enlighten pipeline, with the entire process detailed as follows.
Step 1. Steer: To utilize off-the-shelf LMMs maximally on reasoning, we employ the category of visual trace evidence to steer the reasoning process. We first notify the AI expert with the human-annotated category set $\mathcal{F}_{i}$ for the $i^{th}$ target image to ensure reasoning process focused on forensic trace evidence rather than drifting into unrelated domains. The $\mathcal{F}_{i}$ directly participates in the following two steps.
Step 2. Demonstrate: To further mitigate hallucinations in imperfect models, we leverage human demonstrations as in-context examples, inspired by LMMs’ ability to learn from explanations in context [63]. Considering the trade-off between context length and efficiency, we adaptively select the most relevant demonstrations by filtering a minimum covering set [64] ${\chi}_{i}$ from $\{{\chi}^{+},{\chi}^{-}\}$ for $\mathcal{F}_{i}$. We ensure at least one positive sample and one negative sample are included in ${\chi}_{i}$ via optimizing the following process:
@@
[…]
where $L(\cdot)$ represents the mapping function that extracts the categories covered by a given subset of exemplars. This ensures the selected examples can comprehensively represent the key forensic attributes required for reasoning.
Step 3. Enlighten: The core of ACoTI lies in enlightening the AI expert through a carefully structured prompting scheme. As illustrated in Fig. 3, the proposed ACoTI prompting emulates human reasoning process with few-shot chain-of-thought (CoT) prompting [65]. A unique aspect is explicitly providing the AI with the correct authenticity label in the context, guiding the model toward the appropriate answer space [65]. Finally, this framework is structured as a quadruple: <context, target, human supervisions, CoT prompt>, where human supervision contains $\mathcal{F}_{i}$, $\chi_{i}$ and $auth_{i}$. Detailed prompts are provided in the SM.
@@
In this way, ACoTI provides an effective pipeline for distilling forensic knowledge from strong models into usable reasoning data. Compared with manually collecting long and complex descriptions from human annotators, ACoTI significantly reduces the burden while maintaining data quality. To quantify its effectiveness, we further conducted a two-alternative forced choice (2AFC) study on human preference between the outcomes generated with ACoTI and those without. Six annotators rated on 100 randomly sampled annotated images, and ACoTI demonstrated 99.17% preference, indicating its effectiveness as a cost-efficient approach for extracting forensic knowledge from imperfect models. Therefore, we apply ACoTI to all collected images and organize the resulting data into the predefined reasoning format $rsn$. As shown in Fig. 4(b), we provide an image with corresponding $rsn$. The 47,594 tuples of $\{\bm{\mathtt{I}},auth,rsn\}$ compose the FakeChain dataset, where $\bm{\mathtt{I}}$ and $auth$ denote the image and authenticity label respectively.
IV FakeInstruct: Million-scale Multimodal Forensic Instruction Dataset
@@
Although FakeChain offers corpus-rich reasoning with abundant forensic knowledge, we transform it into a diverse set of visual instructions to better leverage the information embedded in long-form text. As illustrated in Fig. 2(b), we further derive instruction-following data from FakeChain in the visual question answering (VQA) format with both Mistral [66] and GPT-4o [33] in an intermingling manner to enhance linguistic and instructional diversity. We take extra precautions to remove textual contamination allowing models to infer answers without visual information, thereby ensuring the integrity of instruction-based reasoning [67]. As depicted in Fig. 4, the dataset contains 2000K diversified instructions in total, which endows LMMs with flexibility and knowledge in handling extensive forensic scenarios. The details are elaborated as follows.
@@
Part 1. Absolute Authenticity Judgments: As shown in Fig. 4(a), we design diverse question formulations on absolute image authenticity to explicitly implant the concepts of fake and real in LMMs. Throughout this part, we utilize the term fake to denote the AI-generated images, while real for the opposite. While incorporating synonyms such as genuine, fault, and artificial can enhance interaction flexibility with human users, we prioritize a higher level of certainty in objective authenticity assessment in this subset.
@@
Part 2. Holistic Reasoning: As shown in Fig. 4(b), the holistic reasoning imitates human inference on image authenticity, which can foster LMMs’ awareness of the relationships between multi-level visual trace evidence and the overall authenticity. Additionally, the carefully structured cause-and-effect reasoning framework enhances LMMs’ ability to distinguish AI-generated images by activating their latent reasoning capabilities [68].
@@
Part 3. Fine-grained Trace Evidence: To equip LMMs with the ability to seamlessly respond to diverse forensic queries from users, we decompose and reformat the reasoning data in FakeChain into a variety of visual instructions, structured in narrative or multiple-choice question (MCQ) [69] formats. This transformation is facilitated by GPT-4o [33] and Mistral [66], whose complementary strengths ensure a diverse distribution of instructional corpus. As illustrated in Fig. 4(c), the dataset includes authenticity-related VQAs covering what, how, yes-or-no, and imperative question types. These questions pertain to perceptual adjectives (e.g., unnatural, misaligned), questionable visual attributes (e.g., smudge, shape), regions of occurrence (e.g., right eye, apple skin), or contextual knowledge (e.g., shadow in wrong directions). The MCQs comprise one correct answer alongside multiple plausible yet misleading distractors, thereby enhancing models’ ability to discern subtle visual cues while mitigating misconceptions.
@@
Part 4. Extensional Knowledge: Beyond fundamental trace evidence reasoning and forensic analysis, we further incorporate extensional knowledge through open-ended discussions (Fig. 4(d)). Specifically, these discussions cover four major aspects: origins of perceptual misjudgments arising from trace evidence, suggesting improvements to enhance the realism of generation, inferring potential generator characteristics based on forensic evidence, and exploring additional user-driven inquiries related to authenticity.
@@
We herein present a brief analysis of the FakeInstruct dataset. In total, about 2000K visual instructions are included in the FakeInstruct dataset. As shown in Fig. 4(e), quantitatively, the image content type and instruction type are well-balanced. Qualitatively, Fig. 4(c) shows decent diversity and information consistency across instructions.
@@ V FakeScope: Large Multimodal Forensic Model
In this section, we introduce the proposed large multimodal expert model, FakeScope, which leverages state-of-the-art LMMs as its backbone and is further fine-tuned on FakeInstruct to incorporate forensic knowledge. Moreover, we propose a non-trivial probability estimation strategy via token soft-scoring, which could extend the model’s qualitative authenticity judgments to quantifiable probability estimations, without requiring additional explicit training on numerical labels. The details are elaborated as follows.
@@ V-A Architecture and Supervised Instruction Tuning
The architecture of FakeScope follows the LLaVA-style [36] vision-language model (VLM) framework, compromising an image encoder that transforms images into tokenized representations, an LLM for textual interpretation and generation, and a vision-language connector that bridge modalities. As listed in Table I, to validate the universality of FakeInstruct, we incorporate two different model architectures, including the 7B scale LLaVA-v1.5 [36] and mPLUG-Owl2 [37], corresponding to two variants of FakeScope. LMM training typically involves two phases: the cross-modal alignment requiring gigantic-scale data and visual instruction tuning on data of proprietary domain [36]. Given the extensive scale of FakeInstruct, we introduce an auxiliary training stage to perform supervised visual instruction tuning specifically on FakeInstruct. Verified by previous studies [70], adequate tuning on the visual instructions in FakeInstruct is expected to facilitate the model’s visual forensic capabilities while also preserving the general knowledge obtained from earlier-phase training. To optimize efficiency, we freeze the vision encoder and fine-tune only the modality projector and language model.
@@ V-B Probability Estimation via Token Soft-Scoring
TABLE I: Meta structures of LMMs as FakeScope foundation.
[…]
LLaVA-v1.5 (7B) [36] CLIP-ViT-L14 MLP Vicuna-7B
mPLUG-Owl2 [37] CLIP-ViT-L14 Abstractor LLaMA2-7B
TABLE II: Publicly available datasets used for evaluation. The ‡ notion indicates the uncontaminated content of the dataset, and # denotes the unit number.
@@
Multimodal Forensic Capability FakeBench [7] Multiple 3k / 3k ImageNet, DIV2K, COCO 10 Detecting, Interpreting, Reasoning, Fine-grained Analyzing
@@
LMMs typically convey their forensic judgment through hard token selection, i.e., choosing the token with the highest logit when generating responses. Here, “token” refers to the vocabulary-level prediction unit in the model’s output distribution, with each token associated with a logit indicating its selection likelihood. This approach inherently limits their ability to provide numerical probability scores, which are crucial for assessing model confidence. Moreover, general-purpose LMMs are not explicitly trained as classifiers with numerical labels, making direct probability estimation an unpreferred methodology. To address this limitation, FakeScope is trained using binary authenticity judgments and incorporates a token soft-scoring strategy to convert model outputs into quantitative probability estimates. In specific, as depicted in Fig. 5, we define fake and real as anchor tokens and apply softmax normalization over their corresponding logits within a standardized prompting framework:
USER: <Image> What is the authenticity of this image?
ASSISTANT: This image is [auth_token].
To compute probabilities, we extract the logits at the specific position of [auth_token], denoted by $\mathcal{T}=\{\tilde{t}_{j}|_{j=0}^{1}\}=\{\tilde{t}_{fake},\ \tilde{t}_{real}\}$, where $\tilde{t}_{(\cdot)}$ represents case variations of the anchor tokens to enhance robustness. The probability $p_{i}$ of each class is then calculated as follows,
where $z^{(v)}$ represents the logit associated with token $v$. By leveraging token-level probability distributions rather than solely selecting the most probable token, this approach retains nuanced confidence signals that would otherwise be discarded during deterministic text generation. This enables FakeScope to generate quantifiable probability estimates, improving its interpretability as a forensic detector.
@@ VI-A Protocols
Additionally, we perform ablation studies to analyze the impact of different training data components and scales (Sec. VI-E). Detailed experimental setups are provided below.
@@ VI-A1 Implementation Details
Training FakeScope is implemented with PyTorch on eight Nvidia H20-NVLink 98G GPUs, and the evaluations are conducted on one NVIDIA Tesla A100 40G GPU. All training are set to full-schedule mode, with detailed parameter configurations provided in SM. To ensure FakeScope serves as a unified forensic model for AI-generated images, we retain a single checkpoint per base model for all evaluations, ensuring fair comparisons. Furthermore, to prevent data leakage, no images from the testing datasets are reused in the training data. To ensure fairness, no data augmentation strategy is employed throughout the entire training process of FakeScope.
@@ VI-A2 Evaluation Datasets
We evaluate the proposed FakeScope’s multifaceted performance on five commonly used datasets in the area of AI-generated image detection. As listed in Table II, we validate multimodal forensic capabilities on FakeBench [7], covering both binary authenticity and multimodal forensic criteria. To assess generalization to unseen content and generation models, we use AntifakePrompt [71], which contains fake images derived from real Microsoft COCO [74] images that are unseen during training. Additionally, we employ the non-overlapping subset of AIGCDetectBenchmark [72], excluding samples overlapping with ImageNet [62] to avoid contamination and ensure fair generalization evaluation. We further test FakeScope on user-generated images from social networks to support in-the-wild evaluation, using the WildRF [73] and SynthWildX [4] datasets, where authenticity labels are derived from social tags. It is worth noting that the scenarios like deepfakes (e.g. the FF++ dataset [75]) and are treated as out-of-distribution (OOD) cases, given their fundamental difference from fully AI-generated images. Nonetheless, they are included in our generalization evaluation to assess the model’s adaptability.

### v2

@@ III-C1 Task Definition
FakeChain encodes forensic knowledge as a materialized reasoning process that links exhaustive description of the appearance and location of trace evidence to an authenticity conclusion. Compared with free-form descriptions, this structured format has two advantages: (1) it allows for a precise and detailed reflection of the aspects humans focus on when assessing authenticity; and (2) it follows the pattern of human decision-making. By organizing supervision in this way, FakeChain can help to improve human-centered transparency. We emphasize, however, that this materialized reasoning should not be equated with the actual causal basis underlying the model’s authenticity judgment, but rather as a post-hoc explanation.
@@ III-C2 ACoTI Strategy
To obtain large-scale reasoning data cost-effectively, we propose a semi-automatic scheme for extracting forensic knowledge from AI experts with marginal human supervision: Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy. As illustrated in Fig. 3, ACoTI follows a human-machine collaborative framework in which humans participate only in the first two stages by providing partial trace-evidence information. This method follows a Steer-Demonstrate-Enlighten pipeline, detailed as follows.
Step 1. Steer: For the $i$-th candidate image, we use the human-annotated trace-evidence category set $\mathcal{F}_{i}$ to steer the reasoning process. In practice, $\mathcal{F}_{i}$ is carried into the next two stages so that reasoning remains grounded in forensic evidence rather than drifting to irrelevant content.
Step 2. Demonstrate: To further reduce hallucination, we provide human-written demonstrations as in-context examples,  [38]. Considering the trade-off between context length and efficiency, we adaptively select a minimum covering set [5] ${\chi}_{i}$ from the exemplar set $\{{\chi}^{+},{\chi}^{-}\}$ according to $\mathcal{F}{i}$ of the target image. Formally, we solve the following optimization:
@@
$\text{{s.t.}}\quad L(\chi^{\prime})=\mathcal{F}_{i},\ \chi^{\prime}\cap\chi^{+}\neq\emptyset,\ \chi^{\prime}\cap\chi^{-}\neq\emptyset,$
where $L(\cdot)$ maps a subset of exemplars to the union of trace categories they cover. In practice, we select 3 to 5 demonstrations from ${\chi}^{+},{\chi}^{-}$} for each target image to satisfy both coverage and label-diversity constraints. The selection is implemented with a greedy category-covering algorithm with random tie-breaking, yielding a compact yet diverse demonstration set that captures the key forensic attributes while remaining within the context window.
Step 3. Enlighten: The core of ACoTI is a structured prompting scheme that guides the teacher model GPT-4V [54] to produce forensic reasoning. As illustrated in Fig. 3, we adopt few-shot chain-of-thought prompting[70] augmented with human supervision. For each target image, we construct a prompt as a quadruple <context, target, human supervision, CoT prompt> , where human supervision includes the attribute set $\mathcal{F}_{i}$, selected demonstrations $\chi_{i}$, and ground-truth label $auth_{i}$. A unique design choice is to explicitly provide the authenticity label in the context to anchor the reasoning toward the appropriate answer space [70]. For example, when the attribute set $\mathcal{F}_{i}$ = {texture, shadow} and ground-truth label $auth_{i}$ = fake, the model is prompted with task context, target, and human supervision containing attribute guidance and filtered demonstrations. The exact prompting structure is provided in the SM.
@@
In this way, ACoTI provides an effective pipeline for distilling forensic knowledge from strong models into structured reasoning data. Compared with manually collecting descriptions from human annotators, ACoTI substantially reduces the burden while maintaining data quality. To quantify its effectiveness, we conducted a two-alternative forced choice (2AFC) study on human preference between the outcomes generated with ACoTI against vanilla prompting. Six annotators compared 100 randomly sampled images, and ACoTI outputs are preferred in 99.17% of cases, validating it as a cost-efficient strategy for extracting forensic knowledge from imperfect models. Based on this, we apply ACoTI to all collected images and organize the resulting data into the predefined reasoning format rsnrsn. As illustrated in Fig. 3, each image is paired with its rsnrsn, forming the final dataset. In total, the resulting FakeChain dataset consists of 47,594 tuples of $\{\bm{\mathtt{I}},auth,rsn\}$, where $\bm{\mathtt{I}}$ and authauth denote the image and its authenticity label, respectively.
IV FakeInstruct: Multimodal Forensic Instruction Dataset
@@
We transform FakeChain into diverse visual instructions to exploit the information embedded in long-form text. As illustrated in Fig. 2(b), we derive instruction-following data with Mistral [34] and GPT-4o [54] in an intermingling manner to enhance linguistic diversity. We take extra precautions to remove textual contamination, prohibiting models from answering without visual information to ensure visual reasoning [25]. As illustrated in Fig. 4, the dataset contains 2M visual instructions, equipping LMMs to handle a wide range of forensic scenarios. Details are given below.
@@
Part 1. Absolute Authenticity Judgments: As shown in Fig. 4(a), we design diverse questions on absolute image authenticity to explicitly instill the concepts of fake and real in LMMs. We use fake to denote the AI-generated images, while real for the opposite. We prioritize unambiguous authenticity judgments in this subset and thus not include synonyms. In total, this part comprises about 47K instructions.
@@
Part 2. Holistic Reasoning: As shown in Fig. 4(b), holistic reasoning induces inference on image authenticity to help LMMs connect multi-level visual trace evidence with overall image authenticity. The structured cause-and-effect format also helps activate latent reasoning abilities for distinguishing AI-generated images [37].This subset includes 95K instructions.
@@
Part 3. Fine-grained Trace Evidence: To help LMMs handle diverse forensic queries, we decompose FakeChain into visual instructions in both narrative and multiple-choice (MCQ) formats. This transformation is facilitated by GPT-4o [54] and Mistral [34], whose complementary strengths improve instructional diversity. As illustrated in Fig. 4(c), this subset covers authenticity-related questions in what, how, yes-or-no, and imperative forms, targeting perceptual adjectives, questionable visual attributes (e.g., shape), regions of occurrence (e.g., right eye), or contextual knowledge (e.g., shadows in the wrong direction). Each MCQ comprises one correct answer and several plausible misleading distractors, encouraging sensitivity to subtle cues while reducing misconceptions. This part contains about 715K instructions.
@@
Part 4. Extensional Knowledge: Beyond forensic analysis, we further introduce extensional knowledge through open-ended discussions (Fig. 4(d)). These discussions cover four aspects: origins of perceptual misjudgments arising from trace evidence, suggesting improvements to enhance the realism of generation, inferring potential generator characteristics, and other user-driven inquiries related to authenticity. This component consists of more than 1190K visual instructions.
@@
Finally, about 2M visual instructions are included in the FakeInstruct dataset. As shown in Fig. 4(e), the image content type and instruction type are quantitatively well-balanced. Qualitatively, Fig. 4(c) shows decent diversity and information consistency across instructions.
@@ V FakeScope: Large Multimodal Forensic Model
In this section, we introduce the proposed large multimodal expert model, FakeScope, which leverages state-of-the-art LMMs as its backbone and is further fine-tuned on FakeInstruct to incorporate forensic knowledge. Moreover, we propose a non-trivial probability estimation strategy via token soft-scoring, which could extend the model’s qualitative authenticity judgments to quantifiable probability estimations, without requiring additional explicit training on numerical labels. This probability estimation operates purely at inference as designed, and does not rely on any auxiliary trainable modules. We aim to enable FakeScope to support user queries beyond binary detection and to provide post-hoc, human-facing rationales that improve the accessibility of its decisions. The details are elaborated as follows.
@@ V-A Architecture and Supervised Instruction Tuning
FakeScope follows the LLaVA-style vision-language architecture [45]. To verify the generality of FakeInstruct, we instantiate FakeScope with two 7B backbones as variants, including LLaVA-v1.5 [45] and mPLUG-Owl2 [78]. Given the scale of FakeInstruct, we further conduct supervised visual instruction tuning on this dataset to adapt the pretrained LMMs to image forensic while also preserving the general knowledge obtained from earlier-phase trainings [45]. We freeze the vision encoder and fine-tune only the modality projector and language model for efficiency.
@@ V-B Probability Estimation via Token Soft-Scoring
LMMs typically convey their forensic judgment through selecting the token with the highest logit, which inherently limits their ability to provide numerical probability scores. Moreover, since general-purpose LMMs are not explicitly trained as classifiers with numerical labels, direct probability prediction is not a natural formulation. To address this, FakeScope is trained using binary authenticity judgments and incorporates a token soft-scoring strategy that operates purely at inference time to convert model outputs into quantitative probability estimates. As depicted in Fig. 5, we define fake and real as anchor tokens and apply softmax normalization over target logits within a standardized prompting framework:
USER: <Image> What is the authenticity of this image?
ASSISTANT: This image is <auth_token>.
Probabilities are computed from the logits at the position of <auth_token>, where the candidate token sets are defined as $\mathcal{T}=\{\tilde{t}_{j}|_{j=0}^{1}\}=\{\tilde{t}_{fake},\ \tilde{t}_{real}\}$, where $\tilde{t}_{(\cdot)}$ represents case variations of the anchor tokens for robustness. The probability $p_{i}$ of each class is then calculated as follows,
where $z^{(v)}$ represents the logit associated with token $v$. Unlike hard token selection, this approach retains nuanced confidence signals that would otherwise be discarded during deterministic text generation. This enables FakeScope to generate quantifiable probability estimates, improving its explainability as a forensic detector.
@@
TABLE I: Publicly available datasets used for evaluation. The ‡ notion indicates the uncontaminated content of the dataset, and # denotes the unit number.
@@
Multimodal Forensic Capability FakeBench [40] Multiple 3k / 3k ImageNet, DIV2K, COCO 10 Detecting, Interpreting, Reasoning, Fine-grained Analyzing, Forensic Attributing
@@ VI-A Protocols
Additionally, we perform ablation studies to analyze the impact of different training data components, scales, and inference strategies (Sec. VI-E). Detailed experimental setups are provided below.
@@ VI-A1 Implementation Details
FakeScope is trained in PyTorch on eight NVIDIA H20-NVLink 98GB GPUs, and evaluated on a single NVIDIA Tesla A100 40GB GPU. All runs use the full-schedule setting with full-parameter fine-tuning of the language model and modality projector, while the vision encoder is frozen for efficiency and stability. Detailed hyperparameters are provided in the SM. Training is performed for one epoch and completes in approximately 34 hours. For each base model, we retain a single checkpoint for all evaluations to ensure fair comparisons. No data augmentation is applied during training.
@@ VI-A2 Evaluation Datasets
We evaluate FakeScope on five widely used benchmarks for AI-generated image detection. As summarized in Table I, we assess multimodal forensic capabilities on FakeBench [40], which covers both binary authenticity prediction and open-ended forensic criteria. We further extend FakeBench to FakeBenchMCQ with 3,000 close-ended multiple-choice questions on local visual attributes to evaluate fine-grained forensic attribution; details are provided in the SM. To assess generalization to unseen content and generation models, we use AntifakePrompt [8], whose fake images are derived from Microsoft COCO [42] exclusive in training data. Additionally, for AIGCDetectBenchmark [82], we employ the strictly non-overlapping evaluation split by removing all real images from ImageNet [17] and fake images whose underlying real-image provenance traces back to ImageNet. This source-level isolation ensures no content is shared across training and testing. For in-the-wild evaluation, we use WildRF [7] and SynthWildX [12], which consist of user-generated social-media images. Finally, scenarios such as deepfakes (e.g., FF++ [62]) and image super-resolution are treated as out-of-distribution (OOD) cases because they typically modify real images rather than performing end-to-end synthesis; we nevertheless include them in the generalization evaluation.